If Lynn can tyoe a page in P minutes, what piece of the page can she type in 5 minutes?
Answer by stanbon(57984)

If Lynn can type a page in P minutes, what piece of the page can she type in 5 minutes?
x/5 = 1/P
---
x = (5/p) of a page
=====================
Cheers,
Stan H.
=============